Two objects, 1 and 2, are moving along a flat, horizontal track. Object 1 covers a distance 8d in a time of 2t seconds. Object 2 travels a distance of d in a time of t seconds. How does the speed of object 1 (v1) compare to the speed of object 2 (v2)?

(A) 4v1 = v2

(B) v1 = 4v2

(C) v1 = 8v2

(D) v1 = -v2


1
Expert's answer
2021-05-12T17:55:29-0400

Solution.

"l_1=8d;"

"t_1=2t s;"

"l_2=d;"

"t_1=t s;"

"v_1=\dfrac{l_1}{t_1};" "v_1=\dfrac{8d}{2t}=\dfrac{4d}{t};"

"v_2=\dfrac{l_2}{t_2}; v_2=\dfrac{d}{t};"

"v_1=4v_2;"

Answer: "c) v_1=4v_2."